OpenStudy (anonymous):

Solve (x + 6)(x - 4) = -16 solve please {-4, 6} {-4, 2} {-12, 22}

OpenStudy (gorv):

expand

OpenStudy (gorv):

x^2+2x-24=-16

OpenStudy (gorv):

add 16 on both side

OpenStudy (gorv):

x^2+2x-24+16=0 x^2+2x-8=0

OpenStudy (gorv):

now factorise it

OpenStudy (midhun.madhu1987):